# Pioneering Reliable Assessment in Text-to-Image Knowledge Editing: Leveraging a Fine-Grained Dataset and an Innovative Criterion

### v1

@@ Abstract
During pre-training, the Text-to-Image (T2I) diffusion models encode factual knowledge into their parameters. These parameterized facts enable realistic image generation, but they may become obsolete over time, thereby misrepresenting the current state of the world. Knowledge editing techniques aim to update model knowledge in a targeted way. However, facing the dual challenges posed by inadequate editing datasets and unreliable evaluation criterion, the development of T2I knowledge editing encounter difficulties in effectively generalizing injected knowledge. In this work, we design a T2I knowledge editing framework by comprehensively spanning on three phases: First, we curate a dataset CAKE, comprising paraphrase and multi-object test, to enable more fine-grained assessment on knowledge generalization. Second, we propose a novel criterion, adaptive CLIP threshold, to effectively filter out false successful images under the current criterion and achieve reliable editing evaluation. Finally, we introduce MPE, a simple but effective approach for T2I knowledge editing. Instead of tuning parameters, MPE precisely recognizes and edits the outdated part of the conditioning text-prompt to accommodate the up-to-date knowledge. A straightforward implementation of MPE (Based on in-context learning) exhibits better overall performance than previous model editors. We hope these efforts can further promote faithful evaluation of T2I knowledge editing methods. Corresponding author
@@
Hengrui Gu, Kaixiong Zhou, Yili Wang, Ruobing Wang, Xin Wang✧✦ School of Artificial Intelligence, Jilin University Department of Electrical and Computer Engineering, North Carolina State University {guhr22,wangyl21,wangrb22}@mails.jlu.edu.cn kzhou22@ncsu.edu,xinwang@jlu.edu.cn
@@ 1 Introduction
Text-to-image (T2I) diffusion models have gained significant advancements in encoding real-world concepts via bridging the gap between textual descriptions and visual representations Zhang et al. (2023a); Yang et al. (2023); Saharia et al. (2022); Rombach et al. (2022a). By pre-training on a large number of image-caption pairs, these generative models acquire statistical biases on visual concepts such as colors, objects, and personalities. For example, by inputting a text prompt “the CEO of Tesla", the model can generate a portrait of “Elon Musk". While some concepts are ageless, other encoded knowledge facts may become invalid over time (e.g., head of a state) or induce harmful social biases (e.g., implicit gender of CEO). To address this oversight, knowledge editing Bau et al. (2020); Wang et al. (2022); Santurkar et al. (2021); Sinitsin et al. (2020); De Cao et al. (2021); Mitchell et al. (2021); Meng et al. (2022a, b) provides an efficient solution by patching undesirable model outputs without significantly altering the model’s general behavior on unrelated input.
@@
Second, to establish a reliable evaluation strategy for editing, we propose a novel criterion termed adaptive CLIP threshold. Unlike the previous criterion based on classification, this innovative criterion instead focuses on whether the synthesized image is "sufficiently" similar to the target fact. Specifically, this criterion analyzes the CLIP score distribution of ideal synthesized images and utilizes its parameter estimations to calculate a score threshold that quantifies the degree of "sufficiency". Utilizing this score threshold in decision-making can effectively filter out false successful images in editing evaluation scenarios. Our validation experiments supported by Kosmos-2, the state-of-the-art open-source vision-language model Liu et al. (2023); Peng et al. (2023), demonstrate the superiority of the novel criterion, significantly outperforming the current criterion.
@@ 2 Related Work
Text-to-image model editing. Model editing techniques focus on providing stable, targeted updates to model behavior without costly re-training. Related researches have been carried out on a variety of model architectures, such as generative adversarial networks Bau et al. (2020); Wang et al. (2022), image classifiers Santurkar et al. (2021) and LLMs Meng et al. (2022a, b); Mitchell et al. (2021, 2022). (Orgad et al., 2023) formally describes T2I model editing as modifying model’s generative preference for visual concepts (e.g., editing the default color of Roses from Red to Blue). Subsequent studies start to focus on editing factual knowledge in T2I model: Inspiring from language model editing Meng et al. (2022a, b), ReFACT and Diff-quickfix (Arad et al., 2023; Basu et al., 2023) both encode the to-be-edited knowledge into a key-value vector pair, but place it into different model components (MLP or self-attention block). The concurrent work EMCID Xiong et al. (2024) sequentially distributes key-value vector pairs across multiple model layers to enable massive concept editing while preserving generation quality. Unlike above methods, our proposed MPE interprets knowledge editing as prompt editing, where the model remains intact, thereby avoiding catastrophic forgetting.
@@ 3.3 Adaptive CLIP Threshold Criterion
Eq. (3) formulates the new criterion for editing evaluation. To determine the optimal operator $g(\hat{\mu},\hat{\sigma})$ for the knowledge editing task, we conducted a criterion validation experiment. We tested existing editing methods, TIME Orgad et al. (2023) and ReFACT Arad et al. (2023), on the role-editing benchmark RoAD Arad et al. (2023) using several operator choices (e.g., $\hat{\mu}-2\hat{\sigma}$) to make evaluation decisions. Additionally, we selected Kosmos-2 Peng et al. (2023), the best-performing open-source vision-language model for the Celebrity Recognition task Liu et al. (2023), as the pseudo-label generator (see Appendix B for the pseudo-label generation process) [^1] The ability of GPT-4v to perform person identification has been officially prohibited. Thus, Kosmos-2 was chosen.. Fig. 3 presents the Macro-F1 performance of various operator choices and the current classification-based criterion. The results demonstrate that $\hat{\mu}-2\hat{\sigma}$ is the most effective choice among the candidate operators. Furthermore, the adaptive CLIP threshold consistently outperforms the current criterion, indicating its reliability as an evaluation scheme. In later experiments, we set threshold operator $g(\hat{\mu},\hat{\sigma})=\hat{\mu}-2\hat{\sigma}$.
Figure 3: The Macro-F1 of different criterion(threshold operators). Current refers to the current criterion.
@@ 3.4 MPE: A Proposal for Text-to-Image Knowledge Editing
Multiple editing. Real-world scenarios generally involve a vast pool of knowledge updates. To operate in practical applications, MPE adopts a "Memory + Retrieval" strategy (Mitchell et al., 2022; Gu et al., 2023; Song et al., 2024) and introduces an additional Retriever component. Specifically, when receiving multiple edits $\left\{e^{(1)},\dots,e^{(n)}\right\}$, MPE stores all edits in an external memory and embeds their $p_{\mathrm{edit}}^{(i)}$ by the Retriever to construct a retrieval index. Then for each input prompt $p$, the retrieval index returns the key edit $e^{*}$ that is the most relevant (i.e., closest in the embedding space) to $p$, and sends them together to the Router for prompt editing. The complete workflow of MPE is described in Appendix C.
Implementation. The Router and the Editer can be instantiated using various schemes, such as fine-tuning a pre-trained text classification model (Sanh et al., 2019; Devlin et al., 2018) for the Router and a Seq2Seq model (Lewis et al., 2019; Raffel et al., 2020) for the Editer. In this paper, we consider a lightweight, in-context learning-based implementation: We deploy the pre-trained Contriever model (Izacard et al., 2022) locally as the Retriever component and teach the GPT-3.5-turbo API (Ouyang et al., 2022) to work as both the Router and the Editer simultaneously, by our manually designed demonstrations (i.e., input-label pairs). The concrete prompts used are detailed in Appendix D.
@@ 4.3 Multiple Editing Results
We then focus on some specific metrics. The curves in Fig. 6 show that MPE owns remarkable robustness to multiple editing, which potentially compensates its weaknesses in Specificity. Conversely, the robustness of ReFACT and EMCID to multiple editing seems less than ideal: They both experience relatively large performance degradation across all metrics. We hope these results can act as a call to the community to develop more practical and effective editing methods. More quantitative and qualitative results are provided in Appendix E.
@@ Appendix A Statistics and Construction Details of CAKE
Construction Details. Given the powerful text generation capabilities of LLMs Li et al. (2022); Zhang et al. (2023b), we utilize ChatGPT to automatically gather candidate edit prompts $p_{\mathrm{edit}}$ and target prompts $p_{\mathrm{tar}}$ to form fact edits. Specifically, we prompt ChatGPT to:
@@ Appendix B Detailed process of the Criterion Validation Experiments
To and the most effective threshold operator and validate the superiority of our proposed adaptive CLIP threshold, we leverage the Kosmos-2 (Peng et al., 2023) as the pseudo-label generator, enabling the automatic criterion evaluation. Specifically, Kosmos-2 is prompted to conduct celebrity recognition task (Kosmos-2 is the best open-source VLM on this task according to (Liu et al., 2023)).
Following previous settings, we adopt the zero-shot context for Kosmos-2 to execute the visual question answering task. For each synthesised image from existing editing methods, Kosmos-2 is taught to answer the question "Who is the person in this image?" with subsequent four options. One of these options corresponds to the target figure after editing, while the others are randomly selected from a pool of candidate celebrities. A synthesised image is labeled as "successful" only if Kosmos-2 selects the correct option or directly outputs the name of the target figure.
Appendix C Overall Algorithm of MPE
@@
Appendix D Prompts used for In-context Learning
@@
Appendix E More Quantitative and Qualitative Results

### v2

@@ Abstract
During pre-training, the Text-to-Image (T2I) diffusion models encode factual knowledge into their parameters. These parameterized facts enable realistic image generation, but they may become obsolete over time, thereby misrepresenting the current state of the world. Knowledge editing techniques aim to update model knowledge in a targeted way. However, facing the dual challenges posed by inadequate editing datasets and unreliable evaluation criterion, the development of T2I knowledge editing encounter difficulties in effectively generalizing injected knowledge. In this work, we design a T2I knowledge editing framework by comprehensively spanning on three phases: First, we curate a dataset CAKE, comprising paraphrase and multi-object test, to enable more fine-grained assessment on knowledge generalization. Second, we propose a novel criterion, adaptive CLIP threshold, to effectively filter out false successful images under the current criterion and achieve reliable editing evaluation. Finally, we introduce MPE, a simple but effective approach for T2I knowledge editing. Instead of tuning parameters, MPE precisely recognizes and edits the outdated part of the conditioning text-prompt to accommodate the up-to-date knowledge. A straightforward implementation of MPE (Based on in-context learning) exhibits better overall performance than previous model editors. We hope these efforts can further promote faithful evaluation of T2I knowledge editing methods. Our code is available at https://github.com/Hengrui-Gu/T2IKnowledgeEditing. Corresponding author
@@
Hengrui Gu, Kaixiong Zhou, Yili Wang, Ruobing Wang, Xin Wang✧✦ School of Artificial Intelligence, Jilin University IMES, Massachusetts Institute of Technology {guhr22,wangyl21,wangrb22}@mails.jlu.edu.cn kz34@mit.edu,xinwang@jlu.edu.cn
@@ 1 Introduction
Text-to-image (T2I) diffusion models have gained significant advancements in encoding real-world concepts via bridging the gap between textual descriptions and visual representations Zhang et al. (2023a); Yang et al. (2023); Saharia et al. (2022); Rombach et al. (2022a). By pre-training on a large number of image-caption pairs, these generative models acquire statistical biases on visual concepts such as colors, objects, and personalities. For example, by inputting a text prompt “the CEO of Tesla", the model can generate a portrait of “Elon Musk". While some concepts are ageless, other encoded knowledge facts may become invalid over time (e.g., head of a state) or induce harmful social biases (e.g., implicit gender of CEO). To address this oversight, knowledge editing Bau et al. (2020); Wang et al. (2022); Santurkar et al. (2021); Sinitsin et al. (2020); De Cao et al. (2021); Mitchell et al. (2021); Meng et al. (2022a, b); Zhong et al. (2024); Shi et al. (2024); Khandelwal et al. (2024) provides an efficient solution by patching undesirable model outputs without significantly altering the model’s general behavior on unrelated input.
@@
Second, to establish a reliable evaluation strategy for editing, we propose a novel criterion termed adaptive CLIP threshold. Unlike the previous criterion based on classification, this innovative criterion instead focuses on whether the synthesized image is "sufficiently" similar to the target fact. Specifically, this criterion analyzes the CLIP score distribution of ideal synthesized images and utilizes its parameter estimations to calculate a score threshold that quantifies the degree of "sufficiency". Utilizing this score threshold in decision-making can effectively filter out false successful images in editing evaluation scenarios. Our validation experiments supported by Qwen-vl-max, the state-of-the-art open-source vision-language model Liu et al. (2024); Bai et al. (2023) on the celebrity recognition task, demonstrate the superiority of the novel criterion, significantly outperforming the current criterion.
@@ 2 Related Work
Text-to-image model editing. Model editing techniques focus on providing stable, targeted updates to model behavior without costly re-training. Related researches have been carried out on a variety of model architectures, such as generative adversarial networks Bau et al. (2020); Wang et al. (2022), image classifiers Santurkar et al. (2021) and LLMs Meng et al. (2022a, b); Mitchell et al. (2021, 2022). (Orgad et al., 2023) formally describes T2I model editing as modifying model’s generative preference for visual concepts (e.g., editing the default color of Roses from Red to Blue). Subsequent studies start to focus on editing factual knowledge in T2I model: Inspiring from language model editing Meng et al. (2022a, b), ReFACT and Diff-quickfix (Arad et al., 2023; Basu et al., 2023) both encode the to-be-edited knowledge into a key-value vector pair, but place it into different model components (MLP or self-attention block). The concurrent work EMCID Xiong et al. (2024) sequentially distributes key-value vector pairs across multiple model layers to enable massive concept editing while preserving generation quality. Unlike above methods, our proposed MPE interprets knowledge editing as prompt editing, where the model remains intact, thereby avoiding catastrophic forgetting. This
Text-to-image model personalization. The goal of T2I model personalization (Cohen et al., 2022; Gal et al., 2022; Tewel et al., 2023; Sun et al., 2024) is to adapt pre-trained T2I models to user-specific image generation needs, enabling high-quality and diverse synthesis of previously unseen visual concepts. This task is fundamentally different from knowledge editing: It focuses on generating images of novel subjects while preserving its class-specific prior (Ruiz et al., 2023). In contrast, knowledge editing aims to completely rewrite the outdated factual associations within the model, without retaining the originally associated outputs. Since the emergence of visual concepts and factual knowledge updates often occur asynchronously, combining the two techniques together, as outlined in (Arad et al., 2023), is the most efficient approach to developing practical and flexible T2I models.
@@ 3.3 Adaptive CLIP Threshold Criterion
Eq. (3) formulates the new criterion for editing evaluation. To determine the optimal operator $g(\hat{\mu},\hat{\sigma})$ for the knowledge editing task, we conducted a criterion validation experiment. We tested the existing editing method, ReFACT Arad et al. (2023), on the role-editing benchmark RoAD Arad et al. (2023) using several operator choices (e.g., $\hat{\mu}-2\hat{\sigma}$) to make evaluation decisions. Additionally, we selected Qwen-vl-max Bai et al. (2023), the best-performing open-source vision-language model for the Celebrity Recognition task Liu et al. (2024), as the pseudo-label generator (see Appendix B for the pseudo-label generation process) [^1] The ability of GPT-4v to perform person identification has been officially prohibited. Thus, Qwen-vl-max was chosen.. Fig. 3 presents the Macro-F1 performance of various operator choices and the current classification-based criterion. The results demonstrate that $\hat{\mu}-2\hat{\sigma}$ is the most effective choice among the candidate operators [^2] We also conducted experiments using another VLM , Kosmos-2 Peng et al. (2023), for labeling, along with human-annotated labels, which revealed consistent patterns. Detailed results can be found in Appendix C. Furthermore, the adaptive CLIP threshold consistently outperforms the current criterion, indicating its reliability as an evaluation scheme. In later experiments, we set threshold operator $g(\hat{\mu},\hat{\sigma})=\hat{\mu}-2\hat{\sigma}$.
Figure 3: Using Qwen-vl-max as the pseudo-label generator, the Macro-F1 performance across different criterion / threshold operators. Current refers to the current, classification-based criterion.
@@ 3.4 MPE: A Proposal for Text-to-Image Knowledge Editing
Multiple editing. Real-world scenarios generally involve a vast pool of knowledge updates. To operate in practical applications, MPE adopts a "Memory + Retrieval" strategy (Mitchell et al., 2022; Gu et al., 2023; Song et al., 2024) and introduces an additional Retriever component. Specifically, when receiving multiple edits $\left\{e^{(1)},\dots,e^{(n)}\right\}$, MPE stores all edits in an external memory and embeds their $p_{\mathrm{edit}}^{(i)}$ by the Retriever to construct a retrieval index. Then for each input prompt $p$, the retrieval index returns the key edit $e^{*}$ that is the most relevant (i.e., closest in the embedding space) to $p$, and sends them together to the Router for prompt editing. The complete workflow of MPE is described in Appendix D.
Implementation. The Router and the Editer can be instantiated using various schemes, such as fine-tuning a pre-trained text classification model (Sanh et al., 2019; Devlin et al., 2018) for the Router and a Seq2Seq model (Lewis et al., 2019; Raffel et al., 2020) for the Editer. In this paper, we consider a lightweight, in-context learning-based implementation: We deploy the pre-trained Contriever model (Izacard et al., 2022) locally as the Retriever component and teach the GPT-3.5-turbo API (Ouyang et al., 2022) to work as both the Router and the Editer simultaneously, by our manually designed demonstrations (i.e., input-label pairs). The concrete prompts used are detailed in Appendix E.
@@ 4.3 Multiple Editing Results
We then focus on some specific metrics. The curves in Fig. 6 show that MPE owns remarkable robustness to multiple editing, which potentially compensates its weaknesses in Specificity. Conversely, the robustness of ReFACT and EMCID to multiple editing seems less than ideal: They both experience relatively large performance degradation across all metrics. We hope these results can act as a call to the community to develop more practical and effective editing methods. More quantitative and qualitative results are provided in Appendix F.
4.4 Performance Analysis on the Retriever Component
Prompt Type Efficacy Generality KgeMap Compo
#1 100 100 100 100
#10 100 100 98.7 98.3
#25 100 100 97.0 96.0
#50 100 100 96.3 95.3
#All 100 100 96.3 91.7
Table 5: Retrieval Accuracy on the CAKE dataset. The (# num) refers to the size of edit batch. Note that under (# 1), there is no disturbance edit so the accuracy is always 100.
In this paper, we leverage Contriever (Izacard et al., 2022), a pre-trained information retrieval model, as the Retriever component of MPE. To validate its reliability and effectiveness for this task, we report its Retrieval Accuracy–the percentage of edit prompts for which the Retriever correctly identifies the relevant edits–in our multiple editing experiments. From the results in Table 5, we make the following observations:
  • For the easy edit prompt (Efficacy, Generality), Contriever shows no performance degradation under the current editing scale.
  • For the hard edit prompt (KgeMap, Compo), retrieval performance slightly declines as the editing scale increases, indicating that MPE also encounters paraphrase and compositionality challenges. Still, it demonstrates greater robustness compared to parameter-update baselines.
4.5 Time Overhead Analysis of the Adaptive CLIP Threshold
In Section 3.3, we introduce a novel criterion for T2I knowledge editing, referred to as the adaptive CLIP threshold. This approach requires an additional warm-up stage to pre-calculate the decision threshold, which introduces some time overhead.
However, we emphasize that: 1) For each dataset, the warm-up stage is a one-time process, allowing future researchers to bypass this step by using the pre-calculated threshold we provide. 2) Theoretically, this novel criterion reduces evaluation time by half.
Warm-up Stage Time Estimation: We formally estimate the warm-up time $T_{\mathrm{warm}}$. Below are the notations we define for clarity:
  • $t_{\mathrm{gen}}$: The time required for a T2I model to generate an image ($t_{\mathrm{gen}}\approx 7.32s$ for Stable-Diffusion v1.4).
  • $t_{\mathrm{clip}}$: The time required for a CLIP model to compute the CLIP-score once ($t_{\mathrm{clip}}\approx 0.12s$ for Laion’s ViT-G/14).
  • $n_{\mathrm{ideal}}$: The number of ideal images generated to compute the adaptive threshold for each editing entry ($n_{\mathrm{ideal}}=50$ in our experiments).
  • $n$: The number of editing entries in the dataset ($n=1500$ for the CAKE dataset).
The warm-up time can be approximated by the formula $T_{\mathrm{warm}}=(t_{\mathrm{gen}}+t_{\mathrm{clip}})\cdot n_{\mathrm{ideal}}\cdot n$, which gives $T_{\mathrm{warm}}=(7.32s+0.12s)\cdot 50\cdot 1500=155\mathrm{hours}$.
Evaluation Time Estimation: During evaluation, if we generate $n_{\mathrm{seed}}$ images for each editing entry, the current criterion requires approximately $2*t_{\mathrm{clip}}*n_{\mathrm{seed}}*n$ while the adaptive CLIP threshold requires about $t_{\mathrm{clip}}*n_{\mathrm{seed}}*n$. Hence, the novel criterion reduces the evaluation time by half.
@@ Appendix A Statistics and Construction Details of CAKE
Construction Details. Given the powerful text generation capabilities of LLMs Li et al. (2022); Zhang et al. (2023b); Miao et al. (2024); Wang et al. (2024); Shen et al. (2024), we utilize ChatGPT to automatically gather candidate edit prompts $p_{\mathrm{edit}}$ and target prompts $p_{\mathrm{tar}}$ to form fact edits. Specifically, we prompt ChatGPT to:
@@ Appendix B Detailed process of the Criterion Validation Experiments
To validate the superiority of our proposed adaptive CLIP threshold and determine the most suitable, human-aligned threshold operator, we leverage two powerful vision-language models(VLMs), Qwen-vl-max (Bai et al., 2023) and Kosmos-2 (Peng et al., 2023) as the pseudo-label generator, enabling automatic criterion evaluation. Given the excellent performance of current VLMs, it is widely acknowledged that VLM-based automatic validation is fairly reliable and more reproducible than human evaluation (Zhang et al., 2023c).
Specifically, we first apply the T2I knowledge editing method, ReFACT (Arad et al., 2023), to synthesize edited images for the Efficacy prompt corresponding to each role-editing entry in RoAD, the knowledge editing dataset. Next, we utilize predefined prompts to instruct the VLMs to perform the celebrity recognition task on these edited images in a zero-shot manner: Given an edited image as visual input, the VLMs are prompted to answer the question, ‘Who is this person?’ by choosing from four specific role options and a None of the above option. One of the role options corresponds to the target figure after editing, while the others are randomly selected from a pool of candidates of the same gender as the target figure. A synthesized image is labeled as ’successful’ only if the VLMs select the correct option or directly output the name of the target figure.
Appendix C Additional Results from the Criterion Validation Experiments.
Additional results of criterion validation experiments using Kosmos-2 as label generator and human-annotated labels are presented in Fig. 7 and Fig. 8, respectively.
Figure 7: Using Kosmos-2 as the pseudo-label generator, the Macro-F1 performance across different criterion / threshold operators. Current refers to the current, classification-based criterion.
Figure 8: Using human-annotated labels, the Macro-F1 performance across different criterion / threshold operators. Current refers to the current, classification-based criterion.
Appendix D Overall Algorithm of MPE
@@
Appendix E Prompts used for In-context Learning
@@
Appendix F More Quantitative and Qualitative Results